OpenStudy (anonymous):

-5(1-5x)+5(-8x-2)=-4x-8x

OpenStudy (anonymous):

-5+25x-40x-10=-4x-8x -15x+12x=15 -3x=15 x=-5

OpenStudy (anonymous):

here u go :D

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

yeaa! ur welcome~